# Einstein Fields: A Neural Perspective to Computational General Relativity

### v1

@@ A.1.1 Contravariant and covariant components
Loosely speaking, an $n-$dimensional vector $v\in\mathbb{R}^{n}$ can be expanded in its basis as $v=v^{1}e_{1}+v^{2}e_{2}+\ldots+v^{n}e_{n}$, or $v=v^{i}e_{i}$ if we use Einstein sum convention. In general relativity, we write $v=v^{\mu}e_{\mu}$, where Greek indices indicate $4-$ dimensional space-time. Thus, in this non-Euclidean setting, it is necessary to distinguish objects that carry an upper index (contravariant) versus objects that carry a lower index (covariant), since they satisfy different geometric properties and transformation laws.
Definition  1 (Contravariance of vector components): Let $X\subset\mathbb{R}^{n}$ be a coordinate system (frame) that is spanned by a coordinate basis set $\{e_{\mu}\}_{1\leq\mu\leq n}$, i.e., each basis vector can be expressed as $e_{\mu}=\frac{\partial}{\partial x^{\mu}}$. A vector $v\in X$ can be expanded in its coordinate basis as $v=v^{\mu}(x)e_{\mu}:=v^{\mu}(x)\frac{\partial}{\partial x^{\mu}}$. When transforming the vector $v$ to a new coordinate system, spanned by another coordinate basis set $\{\bar{e}_{\nu}\}_{1\leq\nu\leq n}$, i.e., $\bar{e}_{\nu}=\frac{\partial}{\partial\bar{x}^{\nu}}$, one can express the vector components in the new coordinate system $\bar{v}^{\nu}=\bar{v}^{\nu}\big{(}\bar{x}\big{)}$ as
@@
Definition  2 (Covariance of basis set): Let $X\subset\mathbb{R}^{n}$ be a coordinate system (frame) that is spanned by a coordinate basis set $\{e_{\mu}\}_{1\leq\mu\leq n}$, i.e., each basis vector can be expressed as $e_{\mu}=\frac{\partial}{\partial x^{\mu}}$. A vector $v\in X$ can be expanded in its coordinate basis as $v=v^{\mu}(x)e_{\mu}:=v^{\mu}(x)\frac{\partial}{\partial x^{\mu}}$. When transforming the vector $v$ to a new coordinate system, spanned by another coordinate basis set $\{\bar{e}_{\nu}\}_{1\leq\nu\leq n}$, i.e., $\bar{e}_{\nu}=\frac{\partial}{\partial\bar{x}^{\nu}}$, then, the basis set itself transforms as,
@@ A.1.2 Dual space
Definition  2 (Dual vector space): Let ($\mathcal{V}$, $+$, $\cdot$) be a vector space over a field $F$ (e.g., $\mathbb{R},\mathbb{C})$. The (algebraic) dual space ($\mathcal{V}^{*}$, $+$, $\cdot$) is a vector space of linear functionals (maps) $\mathcal{V}^{*}:=\big{\{}v^{*}|v^{*}(v)=c\ \forall v\in\mathcal{V},\forall c\in F\big{\}}$, satisfying:
@@
[…]
@@ A.1.3 Tangent and cotangent spaces
iii) $v(fg)=f(p)v(g)+g(p)v(f)\ \ \forall f,g\in C^{\infty}(\mathcal{M},\mathbb{R})$ (Leibniz product rule)
@@ A.2.2 Operations on tensor fields
I.e., $\alpha_{p}$ is summed-over in the contravariant and covariant indices, and, thus it gets contracted. The resulting tensor is of rank $(r-1,s-1)$
@@ A.2.3 Lie derivative: Generalizing the notion of directional derivatives for tensor fields
In order to compute Equation (42), we need two important concepts from differential geometry called push-forward and pull-back operations. We direct the interested readers to more advanced literature Isham (1999); Lee (2012); Kobayashi & Nomizu (1963)
These three separate tensors fields can be related in the following manner: Firstly, the tensor mapped to the new set of coordinates $\bar{T}^{\mu}_{\nu}(\bar{x})$ can be obtained via Equation(35),
@@
Substituting Equations (43) (44) into the Lie-derivative definition of Equation (42), and $\delta\xi\rightarrow 0$ one finds the following final expression:
@@
Thus, tensors are being compared in the same tangent/cotangent space. Mathematically, for smooth maps777for e.g., dragging of coordinates as in Equation (40) due to flows induced by vector fields. (diffeomorphisms) $\Phi:\mathcal{M}\rightarrow\mathcal{N}$ the pushforward $\Phi_{*}:T_{p}\mathcal{M}\rightarrow T_{\Phi(p)}\mathcal{N}$ pushes vector fields forward from one tangent space of a domain $T_{p}\mathcal{M}$ to the tangent space of another tangent space $T_{\Phi(p)}\mathcal{N}$. The pullback, a dual linear map to pushforward, drags covectors (one-forms) living in cotangent spaces $(\Phi_{*})^{*}\equiv\Phi^{*}:T^{*}_{\Phi(p)}\mathcal{N}\rightarrow T^{*}_{p}\mathcal{M}$ in the reverse direction to the domain. Hence, the contributions from the pushforward on the vector field components and pullback on the covector field components jointly determine the structure of the Lie derivative of a mixed tensor field, as expressed in Equation (45). These operations offer a coherent mathematical framework for transitioning between tangent and cotangent bundles mapped onto other tangent and cotangent bundles via smooth maps, acting appropriately on vector fields and one-forms, respectively.
For any arbitrary rank $(r,s)$ tensor, Equation (45) can be generalized to:
@@ A.3.2 Metric tensor
Geometrically, the metric defined in Equation (49) generalizes the notion of distances and induces a norm ||.$||.||_{p}:T_{p}\mathcal{M}\rightarrow\mathbb{R}$ for generic coordinates such as curvilinear and/or manifolds possessing geometries that are intrinsically non-Euclidean in nature, for e.g., spaces of constant positive sectional curvature $K=1$ (e.g., a 2-sphere $\mathbb{S}^{2}$ embedded in $\mathbb{R}^{3}$), spaces of constant sectional curvature $K=-1$ such as hyperbolic geometry (Bolyai-Lobachevsky spaces $\mathbb{H}^{2}$).
@@
$g_{\mu\nu}=\begin{pmatrix}\langle\phi_{0},\phi_{0}\rangle&\langle\phi_{0},\phi_{1}\rangle&\cdots&\langle\phi_{0},\phi_{n-1}\rangle\\ \langle\phi_{1},\phi_{0}\rangle&\langle\phi_{1},\phi_{1}\rangle&\cdots&\langle\phi_{1},\phi_{n-1}\rangle\\ \vdots&\vdots&\ddots&\vdots\\ \langle\phi_{n-1},\phi_{0}\rangle&\langle\phi_{n-1},\phi_{1}\rangle&\cdots&\langle\phi_{n-1},\phi_{n-1}\rangle\\ \end{pmatrix}\ .$ (51)
@@
the metric tensor amounts to $g_{ij}=\delta_{ij}$. and boils down to Pythagoras’ theorem. A general Riemannian metric prescribes a method to measure the norm of a vector $v$ as $\sqrt{g(v,v)}=||v||$ and also allows for measuring angles between any two vectors $v,w$ at each point $\text{cos}\vartheta=\frac{g(v,w)}{\sqrt{g(v,v)g(w,w)}}$. Like any other tensor, the components of the metric tensor transform under a coordinate change according to Equation (35):
@@
[…]
@@
Unlike Riemannian manifolds, spacetime is actually a pseudo-Riemannian manifold [^9] In general, a pseudo-Riemannian manifold has a signature $(\underbrace{-,...,-}_{m},\underbrace{+,...,+}_{n})$., that is, the metric is not positive definite. Thus, the underlying metric carries a signature (,+,+,+)(-,+,+,+)( - , + , + , + ), meaning, $g_{tt}<0$. Consequently, spacetime is a Lorentzian manifold $\mathscr{M}$, and, forms the basis for electromagnetism and special relativity. The simplest example of a Lorentzian manifold of arbitrary dimension is the Minkowksi metric, which is flat (meaning no curvature):
@@
[…]
@@ A.3.3 Connections & covariant derivative
[…]
where, $[v,w]=\big{(}v^{\mu}\partial_{\mu}w^{\nu}-w^{\mu}\partial_{\mu}v^{\nu}\big{)}\partial_{\nu}$ is the Lie-bracket of vector fields Kobayashi & Nomizu (1963).
@@
[…]
@@
A crucial feature of any connection is that it is not a tensorial quantity. Connections don’t obey the transformation law in Equation (35) under coordinate changes. This can be easily seen through the components of the Christoffel symbols in the coordinate basis:
@@
In case of a nonzero Levi-civita connection, the partial derivatives of an ordinary Lie derivative in Equation(46) is replaced by the covariant derivatives:
@@
The first term advects (drags) the tensor along the flow of $v$, i.e., this is the “naive” directional derivative part. The second and third terms account for how the basis vectors themselves are changing, due to curvature and due to the vector field $v$, respectively. It is easy to show that the three terms lead to pair-wise cancellations between the Christoffel symbols present in the three different covariant derivative terms of Equation(68). Due to this, the whole expression boils down to Equation(46), thus corroborating the connection independence of this derivative operator. Differently put, in the covariant derivatives, the Christoffel symbols introduce extra terms. However, the extra Christoffel terms cancel out between the different contributions (first, second, and third terms). One ends up getting exactly the same final expression for the Lie derivative as if only partial derivatives had been used – this is what is meant by “connection independence”.
@@ A.3.4 Geodesic equation
Geodesics are paths that correspond to the most stationary trajectories (shortest or longest distance) that connect two points $p$ and $q$ on a manifold. Often, we only consider locally distance minimizing curves, and refer to them as geodesics. Geodesics are obtained by solving a calculus of variations problem on the distance metric $\mathcal{L}(\gamma)$ in Equation (54), i.e.,
@@
(or alternatively on $\tau(\sigma)$ in the Lorentzian setting of Equation (57)). Solving the calculus of variations problem boils down to solving the Euler-Lagrange equations, which mathematically is equivalent to the condition
@@
where $\gamma(t)$ is the curve (path) on the manifold, $\dot{\gamma}(t)$ the tangent vector (velocity vector), and $\nabla$ the covariant derivative. Equation (70) intuitively says that the tangent vector is parallel transported along itself – meaning, one is moving without “acceleration” relative to the curved space. Thus, parallel transporting the tangent vector along the curve preserves the tangent vector. In numerical relativity, these corresponds to the equations of motion, i.e. a generalization of the Newton’s acceleration equation $\frac{d^{2}x^{\mu}}{d\tau^{2}}=F^{\mu}/m$. For full derivations, we direct the readers to refer to Carroll et al. (2004); Misner et al. (2017); Poisson (2004).
@@
where, $\tau$ is some affine paramter (typically, chosen to be the proper-time in Equation (57)), $d^{2}x^{\mu}/d\tau^{2}$ is the four-acceleration vector, $dx^{\rho}/d\tau$ is the four-velocity and $\Gamma^{\mu}_{\rho\sigma}$ is the Christoffel symbols as seen in Equation (66)).
Importantly, Equation (70) is the geometric statement of the geodesic equation. It’s coordinate-free, i.e., it’s expressed entirely in terms of geometric objects. Equation (71) is the coordinate version of the same idea. Here, one chooses a coordinate system $x^{\mu}$ on the manifold, and the covariant derivative $\nabla$ acting on a vector becomes the partial derivative plus correction terms involving the Christoffel symbols [^11] Ideally, concepts such as connections, parallel transport and covariant derivatives are metric-independent formulation.

### v2

@@ Acknowledgments
Reproducibility statement
We have made every effort to ensure the reproducibility of our results. The codebase, including training scripts, neural field models used, and data generation and preprocessing pipelines, will be made accessible as a zip file in the supplementary material. All experiments can be reproduced using the instructions provided in the repository’s README.md and How_to_train_EinFields.md, with detailed specifications of hyperparameters, optimizer settings etc.
For our synthetically generated analytic solutions data, we provide the essential configuration yaml files with appropriate parameter values in data_generation/configs and the data generation scripts are within data_generation. Copious example notebooks containing all the validation problems are contained with the folder example_notebooks. The blackhole render scripts and visualization can be found within the bh_render folder.
Software dependencies are specified in a requirements.txt file, and we provide Conda virtual environments for ease of setup, especially with the appropriate CUDA version. All experiments were run on [specify hardware, e.g., 1× NVIDIA A100 GPUs for prototyping and 1 NVIDIA H200 GPUs for the main runs], with a training time ranging from 100 - 2200 seconds depending on Jacobian, Hessian-inclusion in losses and the specific hardware used.
@@ A.1.1 Contravariant and covariant components
Loosely speaking, an $n$-dimensional vector $v\in\mathbb{R}^{n}$ can be expanded in its basis as $v=v^{1}e_{1}+v^{2}e_{2}+\ldots+v^{n}e_{n}$, or $v=v^{i}e_{i}$ if we use Einstein sum convention. In general relativity, we write $v=v^{\mu}e_{\mu}$, where Greek indices indicate 44- dimensional space-time. Thus, in this non-Euclidean setting, it is necessary to distinguish objects that carry an upper index (contravariant) versus objects that carry a lower index (covariant), since they satisfy different geometric properties and transformation laws.
Definition 1 (Contravariance of vector components): Let $X\subset\mathbb{R}^{n}$ be a coordinate system (frame) that is spanned by a coordinate basis set $\{e_{\mu}\}_{1\leq\mu\leq n}$, i.e., each basis vector can be expressed as $e_{\mu}=\frac{\partial}{\partial x^{\mu}}$. A vector $v\in X$ can be expanded in its coordinate basis as $v=v^{\mu}(x)e_{\mu}:=v^{\mu}(x)\frac{\partial}{\partial x^{\mu}}$. When transforming the vector $v$ to a new coordinate system, spanned by another coordinate basis set $\{\bar{e}_{\nu}\}_{1\leq\nu\leq n}$, i.e., $\bar{e}_{\nu}=\frac{\partial}{\partial\bar{x}^{\nu}}$, one can express the vector components in the new coordinate system $\bar{v}^{\nu}=\bar{v}^{\nu}\big(\bar{x}\big)$ as
@@
Definition 2 (Covariance of basis set): Let $X\subset\mathbb{R}^{n}$ be a coordinate system (frame) that is spanned by a coordinate basis set $\{e_{\mu}\}_{1\leq\mu\leq n}$, i.e., each basis vector can be expressed as $e_{\mu}=\frac{\partial}{\partial x^{\mu}}$. A vector $v\in X$ can be expanded in its coordinate basis as $v=v^{\mu}(x)e_{\mu}:=v^{\mu}(x)\frac{\partial}{\partial x^{\mu}}$. When transforming the vector $v$ to a new coordinate system, spanned by another coordinate basis set $\{\bar{e}_{\nu}\}_{1\leq\nu\leq n}$, i.e., $\bar{e}_{\nu}=\frac{\partial}{\partial\bar{x}^{\nu}}$, then, the basis set itself transforms as,
@@ A.1.2 Dual space
Definition 2 (Dual vector space): Let ($\mathcal{V}$, ++, $\cdot$) be a vector space over a field $F$ (e.g., $\mathbb{R},\mathbb{C})$. The (algebraic) dual space ($\mathcal{V}^{*}$, ++, $\cdot$) is a vector space of linear functionals (maps) $\mathcal{V}^{*}:=\big\{v^{*}|v^{*}(v)=c\ \forall v\in\mathcal{V},\forall c\in F\big\}$, satisfying:
@@
for all $v^{*},w^{*}\in\mathcal{V}^{*},\ v,w\in\mathcal{V},\ \alpha,\beta,c\in F$. Elements of the dual space $\mathcal{V}^{*}$ are sometimes referred to as covectors or one-forms Bott & Tu (1982). [^2] A finite dimensional vector space is isomorphic to its double dual, i.e. $\mathcal{V}\cong\big(\mathcal{V}^{*}\big)^{*}$..
@@ A.1.3 Tangent and cotangent spaces
[…]
@@ A.2.2 Operations on tensor fields
I.e., $\alpha_{p}$ is summed-over in the contravariant and covariant indices, and, thus it gets contracted. The resulting tensor is of rank (r1,s1)(r-1,s-1) .
@@ A.2.3 Lie derivative: Generalizing the notion of directional derivatives for tensor fields
In order to compute Eq. (27), we need two important concepts from differential geometry called push-forward and pull-back operations. We direct the interested readers to more advanced literature Isham (1999); Lee (2012); Kobayashi & Nomizu (1963)
These three separate tensors fields can be related in the following manner: Firstly, the tensor mapped to the new set of coordinates $\bar{T}^{\mu}_{\nu}(\bar{x})$ can be obtained via Eq. (20),
@@
Substituting Eqs. (28, 29) into the Lie-derivative definition of Eq. (27), and $\delta\xi\rightarrow 0$ one finds the following final expression:
@@
Thus, tensors are being compared in the same tangent/cotangent space. Mathematically, for smooth maps444for e.g., dragging of coordinates as in Eq. (25) due to flows induced by vector fields. (diffeomorphisms) $\Phi:\mathcal{M}\rightarrow\mathcal{N}$ the pushforward $\Phi_{*}:T_{p}\mathcal{M}\rightarrow T_{\Phi(p)}\mathcal{N}$ pushes vector fields forward from one tangent space of a domain $T_{p}\mathcal{M}$ to the tangent space of another tangent space $T_{\Phi(p)}\mathcal{N}$. The pullback, a dual linear map to pushforward, drags covectors (one-forms) living in cotangent spaces $(\Phi_{*})^{*}\equiv\Phi^{*}:T^{*}_{\Phi(p)}\mathcal{N}\rightarrow T^{*}_{p}\mathcal{M}$ in the reverse direction to the domain. Hence, the contributions from the pushforward on the vector field components and pullback on the covector field components jointly determine the structure of the Lie derivative of a mixed tensor field, as expressed in Eq. (30). These operations offer a coherent mathematical framework for transitioning between tangent and cotangent bundles mapped onto other tangent and cotangent bundles via smooth maps, acting appropriately on vector fields and one-forms, respectively.
For any arbitrary rank $(r,s)$ tensor, Eq. (30) can be generalized to:
@@ A.3.2 Metric tensor
Geometrically, the metric defined in Eq. (34) generalizes the notion of distances and induces a norm $||.||_{p}:T_{p}\mathcal{M}\rightarrow\mathbb{R}$ for generic coordinates such as curvilinear and/or manifolds possessing geometries that are intrinsically non-Euclidean in nature, for e.g., spaces of constant positive sectional curvature $K=1$ (e.g., a 2-sphere $\mathbb{S}^{2}$ embedded in $\mathbb{R}^{3}$), spaces of constant sectional curvature K=1K=-1 such as hyperbolic geometry (Bolyai-Lobachevsky spaces $\mathbb{H}^{2}$).
@@
$g_{\mu\nu}=\begin{pmatrix}\langle\phi_{0},\phi_{0}\rangle&\langle\phi_{0},\phi_{1}\rangle&\cdots&\langle\phi_{0},\phi_{n-1}\rangle\\ \langle\phi_{1},\phi_{0}\rangle&\langle\phi_{1},\phi_{1}\rangle&\cdots&\langle\phi_{1},\phi_{n-1}\rangle\\ \vdots&\vdots&\ddots&\vdots\\ \langle\phi_{n-1},\phi_{0}\rangle&\langle\phi_{n-1},\phi_{1}\rangle&\cdots&\langle\phi_{n-1},\phi_{n-1}\rangle\\ \end{pmatrix}\$. (36)
@@
[…]
@@
$\mathcal{L}(\gamma)=\int_{0}^{1}\left\|\dot{\gamma}(t)\right\|_{\gamma(t)}\,dt=\int_{0}^{1}\sqrt{g_{\mu\nu}(x(t))\,\frac{dx^{\mu}(t)}{dt}\,\frac{dx^{\nu}(t)}{dt}}\,dt\ .$ (39)
@@
Unlike Riemannian manifolds, spacetime is actually a pseudo-Riemannian manifold [^6] In general, a pseudo-Riemannian manifold has a signature $(\underbrace{-,...,-}_{m},\underbrace{+,...,+}_{n})$., that is, the metric is not positive definite. Thus, the underlying metric carries a signature (,+,+,+)(-,+,+,+), meaning, $g_{tt}<0$. Consequently, spacetime is a Lorentzian manifold $\mathscr{M}$, and, forms the basis for electromagnetism and special relativity. The simplest example of a Lorentzian manifold of arbitrary dimension is the Minkowksi metric, which is flat (meaning no curvature):
@@
$\mathcal{\tau}(\sigma)=\int^{\sigma}_{0}\sqrt{-g_{\mu\nu}(x)\,\frac{dx^{\mu}(\sigma^{\prime})}{d\sigma^{\prime}}\,\frac{dx^{\nu}(\sigma^{{}^{\prime}})}{d\sigma^{\prime}}}\,d\sigma^{\prime},$ (42)
@@ A.3.3 Connections & covariant derivative
$\nabla_{v}g=0\ \ \forall v\in\Gamma(T\mathcal{M})\ \ \text{(metricity condition)}$ (45a)
$\nabla_{v}w-\nabla_{w}v=[v,w]\ \ \forall v,w\in\Gamma(T\mathcal{M})\ \ \text{(torsion-free condition)},$ (45b)
where, $v,w]=\big(v^{\mu}\partial_{\mu}w^{\nu}-w^{\mu}\partial_{\mu}v^{\nu}\big)\partial_{\nu}$ is the Lie-bracket of vector fields Kobayashi & Nomizu (1963).
@@
(Parallel transport): Let $\gamma:[0,1]\to\mathcal{M}$ be a smooth curve on the manifold, and let $T$ be a smooth $(r,s)$-rank tensor field defined along the curve $\gamma$. The parallel transport of $T$ along the curve $\gamma(\tau)$ is defined by the condition that its directional covariant derivative along the curve’s tangent vector vanishes:
@@
A crucial feature of any connection is that it is not a tensorial quantity. Connections don’t obey the transformation law in Eq. (20) under coordinate changes. This can be easily seen through the components of the Christoffel symbols in the coordinate basis:
@@
In case of a nonzero Levi-civita connection, the partial derivatives of an ordinary Lie derivative in Eq. (31) is replaced by the covariant derivatives:
@@
The first term advects (drags) the tensor along the flow of $v$, i.e., this is the “naive” directional derivative part. The second and third terms account for how the basis vectors themselves are changing, due to curvature and due to the vector field $v$, respectively. It is easy to show that the three terms lead to pair-wise cancellations between the Christoffel symbols present in the three different covariant derivative terms of Eq. (53). Due to this, the whole expression boils down to Eq. (31), thus corroborating the connection independence of this derivative operator. Differently put, in the covariant derivatives, the Christoffel symbols introduce extra terms. However, the extra Christoffel terms cancel out between the different contributions (first, second, and third terms). One ends up getting exactly the same final expression for the Lie derivative as if only partial derivatives had been used – this is what is meant by “connection independence”.
@@ A.3.4 Geodesic equation
Geodesics are paths that correspond to the most stationary trajectories (shortest or longest distance) that connect two points $p$ and $q$ on a manifold. Often, we only consider locally distance minimizing curves, and refer to them as geodesics. Geodesics are obtained by solving a calculus of variations problem on the distance metric $\mathcal{L}(\gamma)$ in Eq. (39), i.e.,
@@
(or alternatively on $\tau(\sigma)$ in the Lorentzian setting of Eq. (42)). Solving the calculus of variations problem boils down to solving the Euler-Lagrange equations, which mathematically is equivalent to the condition
@@
where $\gamma(t)$ is the curve (path) on the manifold, $\dot{\gamma}(t)$ the tangent vector (velocity vector), and $\nabla$ the covariant derivative. Eq. (55) intuitively says that the tangent vector is parallel transported along itself – meaning, one is moving without “acceleration” relative to the curved space. Thus, parallel transporting the tangent vector along the curve preserves the tangent vector. In numerical relativity, these corresponds to the equations of motion, i.e. a generalization of the Newton’s acceleration equation $\frac{d^{2}x^{\mu}}{d\tau^{2}}=F^{\mu}/m$. For full derivations, we direct the readers to refer to Carroll et al. (2004); Misner et al. (2017); Poisson (2004).
@@
where, $\tau$ is some affine paramter (typically, chosen to be the proper-time in Eq. (42)), $d^{2}x^{\mu}/d\tau^{2}$ is the four-acceleration vector, $dx^{\rho}/d\tau$ is the four-velocity and $\Gamma^{\mu}_{\rho\sigma}$ is the Christoffel symbols as seen in Eq. (51)).
Importantly, Eq. (55) is the geometric statement of the geodesic equation. It’s coordinate-free, i.e., it’s expressed entirely in terms of geometric objects. Eq. (56) is the coordinate version of the same idea. Here, one chooses a coordinate system $x^{\mu}$ on the manifold, and the covariant derivative $\nabla$ acting on a vector becomes the partial derivative plus correction terms involving the Christoffel symbols [^8] Ideally, concepts such as connections, parallel transport and covariant derivatives are metric-independent formulation.